# RouteNet-Gauss: Hardware-Enhanced Network Modeling with Machine Learning

### v2

@@ Abstract
Index Terms:
Network Modeling, Neural Networks, Graph Neural Networks, Spatio-Temporal Graph Neural Networks.
@@ I Introduction
In recent years, there has been significant progress in the development of accurate network models, which have become essential tools for network operators. These models play a pivotal role in facilitating various tasks, including capacity planning [1], topology design [2], and traffic engineering [3]. By providing a controlled and safe environment for testing network configurations, these models eliminate the need to modify production networks directly. This capability enables researchers and operators to explore scenarios that would otherwise be too risky or impractical to test in operational environments.
A widely adopted approach to building these network models involves the use of Discrete Event Simulation (DES) methodologies, such as ns-3 [4] and OMNeT++ [5]. DES-based models are highly valued for their ability to provide packet-level visibility, enabling detailed tracking of individual packets as they traverse a network. This level of granularity has made DES methodologies a standard in a wide range of applications, including protocol debugging, performance estimation, topology design, and SLA assurance. However, these models also face notable challenges (explored more thoroughly in Section II):
@@
    Issue #1 - Computational Complexity: DES-based simulators are computationally intensive, as they simulate each individual event in a network (e.g., packet transmission), which can become prohibitive in scenarios involving millions or billions of packets per second [6].
@@
    Issue #2 - Potential Inaccuracies: DES-based simulators often rely on idealized assumptions and scenarios, which may not fully capture the complexities of real-world networks [7, 8]. Furthermore, the lack of access to proprietary details, such as the precise configurations of commercial hardware devices (e.g., queue sizes), potentially limits their accuracy.
Significant research efforts have focused on reducing the computational cost of DES-based simulations. For example, novel DES designs like DONS [9] have introduced parallelization techniques, achieving remarkable speedups - up to 65 times faster than the standard OMNeT++. Similarly, hybrid approaches such as DeepQueueNet (DQN) [10] leverage Machine Learning (ML) models to accelerate specific components of the simulation pipeline, offering comparable efficiency improvements. These advances represent significant steps toward improving simulation performance. However, these methods do not address Issue #2, and are therefore constrained by the accuracy ceiling of DES itself. This includes ML models trained on simulated data [9, 6, 10, 11].
@@
This approach, however, involves a trade-off. To ensure computational efficiency when modeling high-speed traffic, our method operates at flow-level granularity rather than packet-level granularity. While many existing solutions rely on packet-level analysis, flow-level granularity can effectively meet the requirements of various network operation tasks, such as capacity planning, QoS assurance, topology design, traffic engineering, and performance estimation [6, 12, 11]. This simplification is particularly suited for scenarios where aggregated flow-level information provides sufficient detail for practical applications.
@@ II Motivation
Despite its strengths, the event-driven design of DES introduces certain limitations, particularly when it comes to parallelization. Simulators like ns-3 [4] and OMNeT++ [5] struggle with efficiently parallelizing event processing due to the inherent sequential nature of identifying and processing events. While these tools can divide network scenarios into smaller segments to process independently, events that span across multiple segments require inter-process communication. This overhead can significantly reduce overall performance and scalability [9].
To address these limitations, current state-of-the-art solutions focus on enhancing DES-based simulators with parallelization techniques (e.g., DONS [9]) or by incorporating ML models to improve performance (e.g., DQN [10]). While these approaches have made notable progress, DES-based simulations remain computationally expensive. Even with parallelization, the huge volume of events in large-scale network scenarios can make processing impractical. For instance, a single 1 Gbps link may process 250,000 500-byte packets per second, illustrating the immense computational demand of DES-based methods.
@@ II-A DES’s elevated computational cost
To illustrate the high computational cost of DES, we refer to Figure 2, showing the relationship between simulation time and the number of packets in a network scenario. This analysis is performed by using the same network setup while incrementally increasing the traffic volume. The dashed lines in the figure provide an approximation of the expected simulation times as the number of packets grows. Additionally, the dotted vertical line marks the number of packets transmitted by a 10 Gbps link operating at 80% load for one second. This bandwidth reflects the minimum link capacities typically found in modern data centers [13, 14]. The 80% load represents a realistic peak for link utilization, often observed at an average network workload of 50% [15].
@@ II-B Inaccuracy of DES when compared to reality
TABLE I: Differences in average packet delay per-flow per windows between OMNeT++ [5] and the testbed.
@@
In short, network simulators face challenges in accurately modeling real-world networks. These limitations inevitably extend to current state-of-the-art ML models that rely on simulated data for training [9, 6, 10, 11], as the quality of an ML model is inherently tied to the quality and representativeness of the data it is trained on. Thus, any inaccuracies or gaps in the simulator’s ability to replicate network behavior can degrade the performance of these models when applied to real-world scenarios.
@@ III-A RouteNet-Gauss’s network decomposition
RouteNet-G takes as input a network scenario, including the traffic, routing, and network configuration, and outputs per-flow performance metrics. This includes critical metrics such as delay, packet loss, and queue occupancy. By training on high-fidelity testbed data, RouteNet-G is designed to model these scenarios accurately while maintaining the flexibility to generalize to unseen network topologies and configurations.
A straightforward approach —such as to model the entire network as a single neural network trained end-to-end— might seem appealing but introduces critical challenges. Specifically, such an architecture would be heavily biased toward the specific topologies and scenarios used during training, making it difficult to generalize to unseen or larger networks. To overcome this, RouteNet-G adopts a divide-and-conquer strategy: it models the individual elements in the network (i.e., queues, links, devices, and flows) and, more importantly, learns how these elements interact.
@@
RouteNet-G’s architecture is inspired by Message Passing Neural Networks (MPNN)[16], a class of GNNs [17]. MPNNs are specifically designed to handle graph-structured data and have demonstrated strong generalization capabilities across unseen graphs. RouteNet-G builds upon this foundation, extending and refining the architecture introduced in RouteNet-Fermi [11] to incorporate the additional complexities of dynamic routing, flow-level granularity, and real-world testbed data.
@@ III-C RouteNet-Gauss’s architecture
$\mathcal{F},\mathcal{L},\mathcal{Q},\mathcal{N},\boldsymbol{x}_{f},\boldsymbol{x}_{l},\boldsymbol{x}_{q}$$\hat{\boldsymbol{y}}_{f}$$q\in\mathcal{Q}$$\boldsymbol{h}^{(0)}_{q}\leftarrow E_{q}(\boldsymbol{x_{q}})$$d\in\mathcal{N}$
[…]
@@
[…]
@@
TABLE II: Extracted features of each element in the scenario.
We now proceed to explain RouteNet-G’s architecture, which follows the same three-part structure as MPNNs: (1) encoding of the internal states for each element, (2) the Message Passing (MP) phase, where the internal states are updated with the information from neighboring elements, and (3) the readout phase, where the outputs of the MP phase are used to obtain the desired performance metrics. Algorithm 1 represents RouteNet-G’s architecture in pseudocode, while the MP and readout phases are further detailed in Algorithm 2.
First, the encoding of the internal states is shown in Algorithm 1, lines 1-2. The encoding takes the relevant input features for each network element, listed in Table II, and encodes them using a Multi-Layer Perceptron (MLP). In the case of the initial states for network devices, $h^{0}_{d}$, the update function $U_{D}$ from the MP algorithm (to be discussed later) is used instead. The loop present in line 5 represents the temporal component of the algorithm. In it, the MP algorithm is executed for each window (line 3).
The MP phase is presented in Algorithm 2, lines 2-12. It is an iterative process, in which the internal states of each element are updated according to their interactions with each other. The outer loop represents the number of times the states of the network elements are updated. In its body, the “building blocks” NNs introduced back in Section III-A are executed, here referred to as update functions $U$. Each function takes as input the relevant network elements and updates the current one. Note again that the same $U$ functions are applied for all network elements and across all windows, as otherwise, the model would not adapt to unseen topologies. All of these functions are modeled through the use of Gated Recurrent Unit (GRU) [18] cells. Finally, in the readout phase (lines 12-13), the readout function $R$ takes as input the final flow states and returns the predicted performance metrics. Function $R$ is implemented through an MLP.
@@ V-A Evaluation of RouteNet-G’s temporal cost
Figure 5: RouteNet-G’s inference cost against DES.
First, we compare RouteNet-G’s performance in evaluating network scenarios against current state-of-the-art methods. We specifically assess the inference time of the trained RouteNet-G model relative to other alternatives, including OMNeT++ [5], DONS [9], and DQN [10]. Both RouteNet-G and the baseline methods are executed in a single thread, except for DQN, which utilizes 4 GPUs. The scenarios evaluated correspond to those illustrated in Figure 2, back in Section II. We analyze the inference time required by each method to process these scenarios and generate its corresponding predictions.
@@ V-B Evaluation of RouteNet-G’s accuracy
Next, we evaluate RouteNet-G’s accuracy when predicting flow performance metrics compared to OMNeT++ [5]. As discussed earlier, current state-of-the-art solutions are either DES-based simulators (e.g., DONS, Parsimon) or ML models trained with samples generated through DES (e.g., DQN). Hence, using OMNeT++, an established DES-based simulator, acts as an upper bound for the accuracy achievable by the current network modeling solutions based on DES.
@@
    Real-World Packet Traces (RWPT): generated flows follow the traffic distribution from real-world traffic traces in the MAWI Working Group Traffic Archive [20].
For each distribution 1823, 700, and 2190 samples were generated, respectively, and then split into training, validation, and test partitions following a 75/15/10 ratio. Partitioning the datasets is common practice to later ensure that the model has not over-fitted to the training data. Each of the samples in these datasets represents a network scenario, consisting of a combination of network topology, input traffic, routing configuration, and a collection of per-flow metrics. Each dataset was generated across 11 distinct topologies, ranging from 5 to 8 nodes. Illustrations of these are recollected in Appendix B. Flow routings span paths traveling from 3 to 5 routers, resulting in an overall range of 12 to 28 hops in the link layer. Moreover, in each network scenario, the number of generated packets varies from 450 thousand to 16 million.

### v3

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
@@ I Introduction
In recent years, there has been significant progress in the development of accurate network models, which have become essential tools for network operators. These models play a pivotal role in facilitating various tasks, including capacity planning [network-planning-21], topology design [topology-discovery18], and traffic engineering [auto-dc-cong-control-18]. By providing a controlled and safe environment for testing network configurations, these models eliminate the need to modify production networks directly. This capability enables researchers and operators to explore scenarios that would otherwise be too risky or impractical to test in operational environments.
A widely adopted approach to building these network models involves the use of Discrete Event Simulation (DES) methodologies, such as ns-3 [Riley2010] and OMNeT++ [Varga2019]. DES-based models are highly valued for their ability to provide packet-level visibility, enabling detailed tracking of individual packets as they traverse a network. This level of granularity has made DES methodologies a standard in a wide range of applications, including protocol debugging, performance estimation, topology design, and SLA assurance. However, these models also face notable challenges (explored more thoroughly in Section II):
@@
    Issue #1 - Computational Complexity: DES-based simulators are computationally intensive, as they simulate each individual event in a network (e.g., packet transmission), which can become prohibitive in scenarios involving millions or billions of packets per second [10.1145/3452296.3472926].
@@
    Issue #2 - Potential Inaccuracies: DES-based simulators often rely on idealized assumptions and scenarios, which may not fully capture the complexities of real-world networks [216073, 10.1145/3508026]. Furthermore, the lack of access to proprietary details, such as the precise configurations of commercial hardware devices (e.g., queue sizes), potentially limits their accuracy.
Significant research efforts have focused on reducing the computational cost of DES-based simulations. For example, novel DES designs like DONS [10.1145/3603269.3604844] have introduced parallelization techniques, achieving remarkable speedups - up to 65 times faster than the standard OMNeT++. Similarly, hybrid approaches such as DeepQueueNet (DQN) [10.1145/3544216.3544248] leverage Machine Learning (ML) models to accelerate specific components of the simulation pipeline, offering comparable efficiency improvements. These advances represent significant steps toward improving simulation performance. However, these methods do not address Issue #2, and are therefore constrained by the accuracy ceiling of DES itself. This includes ML models trained on simulated data [10.1145/3603269.3604844, 10.1145/3452296.3472926, 10.1145/3544216.3544248, ferriolgalmés2022routenetfermi].
@@
This approach, however, involves a trade-off. To ensure computational efficiency when modeling high-speed traffic, our method operates at flow-level granularity rather than packet-level granularity. While many existing solutions rely on packet-level analysis, flow-level granularity can effectively meet the requirements of various network operation tasks, such as capacity planning, QoS assurance, topology design, traffic engineering, and performance estimation [10.1145/3452296.3472926, ferriol2022routeneterlang, ferriolgalmés2022routenetfermi]. This simplification is particularly suited for scenarios where aggregated flow-level information provides sufficient detail for practical applications.
@@ II Motivation
Despite its strengths, the event-driven design of DES introduces certain limitations, particularly when it comes to parallelization. Simulators like ns-3 [Riley2010] and OMNeT++ [Varga2019] struggle with efficiently parallelizing event processing due to the inherent sequential nature of identifying and processing events. While these tools can divide network scenarios into smaller segments to process independently, events that span across multiple segments require inter-process communication. This overhead can significantly reduce overall performance and scalability [10.1145/3603269.3604844].
To address these limitations, current state-of-the-art solutions focus on enhancing DES-based simulators with parallelization techniques (e.g., DONS [10.1145/3603269.3604844]) or by incorporating ML models to improve performance (e.g., DQN [10.1145/3544216.3544248]). While these approaches have made notable progress, DES-based simulations remain computationally expensive. Even with parallelization, the huge volume of events in large-scale network scenarios can make processing impractical. For instance, a single 1 Gbps link may process 250,000 500-byte packets per second, illustrating the immense computational demand of DES-based methods.
@@ II-A DES’s elevated computational cost
To illustrate the high computational cost of DES, we refer to Figure 2, showing the relationship between simulation time and the number of packets in a network scenario. This analysis is performed by using the same network setup while incrementally increasing the traffic volume. The dashed lines in the figure provide an approximation of the expected simulation times as the number of packets grows. Additionally, the dotted vertical line marks the number of packets transmitted by a 10 Gbps link operating at 80% load for one second. This bandwidth reflects the minimum link capacities typically found in modern data centers [10.1145/2829988.2787472, Vahdat_2015]. The 80% load represents a realistic peak for link utilization, often observed at an average network workload of 50% [10.1145/3341302.3342085].
@@ II-B Inaccuracy of DES when compared to reality
Table I: Differences in average packet delay per-flow per windows between OMNeT++ [Varga2019] and the testbed.
@@
In short, network simulators face challenges in accurately modeling real-world networks. These limitations inevitably extend to current state-of-the-art ML models that rely on simulated data for training [10.1145/3603269.3604844, 10.1145/3452296.3472926, 10.1145/3544216.3544248, ferriolgalmés2022routenetfermi], as the quality of an ML model is inherently tied to the quality and representativeness of the data it is trained on. Thus, any inaccuracies or gaps in the simulator’s ability to replicate network behavior can degrade the performance of these models when applied to real-world scenarios.
@@ III-A RouteNet-Gauss’s network decomposition
RouteNet-G takes as input a network scenario, including the traffic, routing, and network configuration, and outputs per-flow performance metrics. This includes critical metrics such as delay, packet loss, and queue occupancy.
RouteNet-G’s architecture works at a flow level. However, flows can be aggregated from individual flows to origin-destination (OD) flows —aggregating all flows that share the same routing path into one— depending on the desired level of granularity. Aggregating flows into OD-flows increases the model efficiency by reducing the number of elements to consider, at the cost of obtaining aggregated predictions. Furthermore, aggregation can be specifically useful when considering “mice flows”, flows with small payloads and flow completion times which form the overwhelming majority of traffic in modern networks [kandula2009nature, jurkiewicz2021flow]. While individually they may be too small for their impact to be noticeable, their aggregate behavior can be more easily captured. In this paper, we will consider OD-flows, as they provide the level of granularity required for the use cases mentioned in Section I.
By training on high-fidelity testbed data, RouteNet-G is designed to model these scenarios accurately while maintaining the flexibility to generalize to unseen network topologies and configurations. A straightforward approach —such as to model the entire network as a single neural network trained end-to-end— might seem appealing but introduces critical challenges. Specifically, such an architecture would be heavily biased toward the specific topologies and scenarios used during training, making it difficult to generalize to unseen or larger networks. To overcome this, RouteNet-G adopts a divide-and-conquer strategy: it models the individual elements in the network (i.e., queues, links, devices, and flows) and, more importantly, learns how these elements interact.
@@
RouteNet-G’s architecture is inspired by Message Passing Neural Networks (MPNN)[10.5555/3305381.3305512], a class of GNNs [scarselli2008graph]. MPNNs are specifically designed to handle graph-structured data and have demonstrated strong generalization capabilities across unseen graphs. RouteNet-G builds upon this foundation, extending and refining the architecture introduced in RouteNet-Fermi [ferriolgalmés2022routenetfermi] to incorporate the additional complexities of dynamic routing, flow-level granularity, and real-world testbed data.
@@ III-C RouteNet-Gauss’s architecture
1: $\mathcal{F},\mathcal{L},\mathcal{Q},\mathcal{N},\boldsymbol{x}_{f},\boldsymbol{x}_{l},\boldsymbol{x}_{q}$
2: $\hat{\boldsymbol{y}}_{f}$
3: for all $q\in\mathcal{Q}$ do $\boldsymbol{h}^{(0)}_{q}\leftarrow E_{q}(\boldsymbol{x_{q}})$
4: for all $d\in\mathcal{N}$ do
5:   $M^{0}_{d}\leftarrow\sum_{q\in d}\boldsymbol{h}^{0}_{q}$
6:   $\boldsymbol{h}^{0}_{d}\leftarrow U_{D}(M^{0}_{d};\boldsymbol{0})$
7: for $t\leftarrow 0$ to T1T-1 do
8:   for all $f\in\mathcal{F}$ do $\boldsymbol{h}^{(t)}_{f}\leftarrow E_{f}(\boldsymbol{x^{(t)}_{f}})$
9:   for all $l\in\mathcal{L}$ do $\boldsymbol{h}^{(t)}_{f}\leftarrow E_{l}(\boldsymbol{x^{(t)}_{l}})$
10:   $\boldsymbol{\hat{y}}^{(t)}_{f},\{\boldsymbol{h}^{(t+1)}_{q}|\forall q\in\mathcal{Q}\},\{\boldsymbol{h}^{(t+1)}_{d}|\forall d\in\mathcal{D}\}\leftarrow MP\big(\{\boldsymbol{h}^{(t)}_{f}|\forall f\in\mathcal{F}\},\{\boldsymbol{h}^{(t)}_{l}|\forall l\in\mathcal{L}\},\{\boldsymbol{h}^{(t)}_{q}|\forall q\in\mathcal{Q}\},\{\boldsymbol{h}^{(t)}_{d}|\forall d\in\mathcal{D}\}\big)$
@@
1: $\boldsymbol{h}^{0}_{f},\boldsymbol{h}^{0}_{l},\boldsymbol{h}^{0}_{q},\boldsymbol{h}^{0}_{d};\forall f\in\mathcal{F},l\in\mathcal{L},q\in\mathcal{Q},d\in\mathcal{N}$
2: $\hat{\boldsymbol{y}}_{f},\boldsymbol{h}^{T}_{q},\boldsymbol{h}^{T}_{d};\forall f\in\mathcal{F},q\in\mathcal{Q},d\in\mathcal{N}$
3: for $t\leftarrow 1$ to $T$ do $\triangleright$ Message Passing Phase
4:   for all $f\in\mathcal{F}$ do $\triangleright$ Message Passing on Flows
5:    $pos\leftarrow 0$ $\triangleright$ Curr. position in path
6:    $\boldsymbol{h}^{t}_{f}\leftarrow\boldsymbol{h}^{t-1}_{f}$
7:    for all $(l,q)\in f$ do
8:      $\widetilde{m}^{t}_{f,pos},\boldsymbol{h}^{t}_{f}\leftarrow U_{F}(\boldsymbol{h}^{t-1}_{l}||\boldsymbol{h}^{t-1}_{q};\boldsymbol{h}^{t}_{f})$
9:      $pos\leftarrow pos+1$
10:   for all $q\in\mathcal{Q}$ do $\triangleright$ Message Passing on Queues
11:    $M^{t}_{q}\leftarrow\boldsymbol{h}^{(t-1)}_{\hat{Q}_{d}(q)}\mathop{\parallel}\displaylimits\sum_{(f,pos)\in\hat{Q}_{f}(q)}{\widetilde{m}^{t}_{(f,pos)}}$
12:    $\boldsymbol{h}^{t}_{q},\emptyset\leftarrow U_{Q}(M^{t}_{q};\boldsymbol{h}^{t-1}_{q})$
13:    $\widetilde{m}^{t}_{q}\leftarrow\boldsymbol{h}^{t}_{q}$
14:   for all $d\in\mathcal{D}$ do $\triangleright$ Message Passing on Devices
15:    $M^{t}_{d}\leftarrow\sum_{q\in d}{\widetilde{m}^{t}_{q}}$
16:    $\boldsymbol{h}^{t}_{d},\emptyset\leftarrow U_{D}(M^{t}_{d};\boldsymbol{h}^{t-1}_{d})$
17:   for all $l\in\mathcal{L}$ do $\triangleright$ Message Passing on Links
18:    $M^{t}_{l}\leftarrow\widetilde{m}^{t}_{\hat{Q}_{q}(l)}$
19:    $\boldsymbol{h}^{t}_{l},\emptyset\leftarrow U_{L}(M^{t}_{l};\boldsymbol{h}^{t-1}_{l})$
20:    $\widetilde{m}^{t}_{l}\leftarrow\boldsymbol{h}^{t}_{l}$
21: for all $f\in\mathcal{F}$ do $\triangleright$ Readout Phase
22:   $\boldsymbol{\hat{y}}_{f}\leftarrow\sum_{\widetilde{m}^{t}_{f,pos}\in\widetilde{m}^{t}_{f}}R(\widetilde{m}^{t}_{f,pos})$
Symbol Definition / Description
$\mathcal{F}$ Set of flows in a network scenario
$\mathcal{L}$ Set of links in a network scenario
$\mathcal{Q}$ Set of queues in a network scenario
$\mathcal{N}$ Set of network devices (nodes) in a network scenario
$E$ Initial encoding function
$U$ Update function of the MP phase
$R$ Readout function
$T$ Maximum iterations of the MP phase; a hyperparameter
$x_{q}$ Input features of queue $q$
$x^{(t)}_{f}$ Input features of flow $f$ during window $t$
$h_{q}^{0}$ Initial encoding for queue $q$
$h_{q}^{t}$ Internal encoding for queue $q$ after iteration $t$ of the MP phase
$h_{q}^{(t)}$ Internal encoding for queue $q$ after the window $t$ of scenario
$\widetilde{m}^{t}_{q}$ Message generated by queue $q$ to be used by its neighbors at iteration $t$
$\widetilde{m}^{t}_{f,pos}$ Message generated by flow $f$ at position pospos of its routing path to be used by its neighbors at iteration $t$
$M_{q}^{t}$ Aggregation of messages for queue $q$ at iteration $t$ of the MP phase
$\hat{Q}_{d}(q)$ Mapping function which returns the device the queue $q$ belongs to
$\hat{Q}_{f}(q)$ Mapping function which returns the flows which include the queue $q$ in its routing path
$\hat{Q}_{q}(l)$ Mapping function which returns the exit queue of link $l$
$\emptyset$ Discarded output from an update function
$\hat{y}^{(t)}_{f}$ Predicted performance metrics of flow $f$ at window $t$
Table II: Table of notations
@@
Table III: Extracted features of each element in the scenario.
We now proceed to explain RouteNet-G’s architecture, which follows the same three-part structure as MPNNs: (1) encoding of the internal states for each element, (2) the Message Passing (MP) phase, where the internal states are updated with the information from neighboring elements, and (3) the readout phase, where the outputs of the MP phase are used to obtain the desired performance metrics. Algorithm 1 represents RouteNet-G’s architecture in pseudocode, while the MP and readout phases are further detailed in Algorithm 2. The symbols used are also described in Table II.
First, the encoding of the internal states is shown in Algorithm 1, lines 3-9. The encoding takes the relevant input features for each network element, listed in Table III, and encodes them using a Multi-Layer Perceptron (MLP). In the case of the initial states for network devices, $h^{0}_{d}$, the update function $U_{D}$ from the MP algorithm (to be discussed later) is used instead. The loop present in line 5 represents the temporal component of the algorithm. In it, the MP algorithm is executed for each window (line 10).
The MP phase is presented in Algorithm 2, lines 3-20. It is an iterative process, in which the internal states of each element are updated according to their interactions with each other. The outer loop represents the number of times the states of the network elements are updated. In its body, the “building blocks” NNs introduced back in Section III-A are executed, here referred to as update functions $U$. Each function takes as input the relevant network elements and updates the current one. Note again that the same $U$ functions are applied for all network elements and across all windows, as otherwise, the model would not adapt to unseen topologies. All of these functions are modeled through the use of Gated Recurrent Unit (GRU) [cho2014learning] cells. Finally, in the readout phase (lines 21-22), the readout function $R$ takes as input the final flow states and returns the predicted performance metrics. Prediction is done for each queue in the flow’s routing path, and then added. Function $R$ is implemented through an MLP.
@@ V-A Evaluation of RouteNet-G’s temporal cost
Figure 5: RouteNet-G’s inference cost against DES. Dashed lines indicate extrapolation from measured values, while the vertical dotted line indicates the expected number of packets seen per second in a modern datacenter link (10 Gbps link at 80% load).
First, we compare RouteNet-G’s performance in evaluating network scenarios against current state-of-the-art methods. We specifically assess the inference time of the trained RouteNet-G model relative to other alternatives, including OMNeT++ [Varga2019], DONS [10.1145/3603269.3604844], and DQN [10.1145/3544216.3544248]. Both RouteNet-G and the baseline methods are executed in a single thread, except for DQN, which utilizes 4 GPUs. The scenarios evaluated correspond to those illustrated in Figure 2, back in Section II. We analyze the inference time required by each method to process these scenarios and generate its corresponding predictions.
@@ V-B Evaluation of RouteNet-G’s accuracy
Next, we evaluate RouteNet-G’s accuracy when predicting flow performance metrics compared to OMNeT++ [Varga2019]. As discussed earlier, current state-of-the-art solutions are either DES-based simulators (e.g., DONS, Parsimon) or ML models trained with samples generated through DES (e.g., DQN). Hence, using OMNeT++, an established DES-based simulator, acts as an upper bound for the accuracy achievable by the current network modeling solutions based on DES. We also compare against our implementation of RouteNet-F, based on the public one available [^2] https://github.com/BNN-UPC/RouteNet-Fermi, and adapted and trained with the testbed datasets.
Traffic Distribution TREX-S TREX-MB RWPT
Number Training $1376$ $541$ $1682$
of Validation $274$ $108$ $336$
network Test $173$ $51$ $172$
scenarios Total $1823$ $700$ $2190$
Number Training $12.64\times 10^{6}$ $8.75\times 10^{6}$ $12.47\times 10^{6}$
of Validation $2.56\times 10^{6}$ $1.67\times 10^{6}$ $2.56\times 10^{6}$
valid Test $1.67\times 10^{6}$ $8.75\times 10^{5}$ $1.24\times 10^{6}$
predictions Total $6.87\times 10^{6} 11$$1.30\times 10^{6}$ $16.27\times 10^{6}$
Table IV: Summary of datasets generated. Valid predictions are defined as the product between the number of flows and extracted windows per scenario.
@@
    Real-World Packet Traces (RWPT): generated flows follow the traffic distribution from real-world traffic traces in the MAWI Working Group Traffic Archive [10.5555/1267724.1267775].
The recorded network scenarios are split into training, validation, and test partitions. Partitioning the datasets is common practice to later ensure that the model has not over-fitted to the training data. Each of the samples in these datasets represents a network scenario, consisting of a combination of network topology, input traffic, routing configuration, and a collection of per-flow metrics. Each dataset was generated across 11 distinct topologies, ranging from 5 to 8 nodes. Illustrations of these are recollected in Appendix B. Flow routings span paths traveling from 3 to 5 routers, resulting in an overall range of 12 to 28 hops in the link layer. Moreover, in each network scenario, the number of generated packets varies from 450 thousand to 16 million. The number of scenarios per type and partition is indicated in Table IV.